Sabas Company has 20,000 shares of $100 par, 2% cumulative preferred stock and 100,000 shares of $50 par common stock. The following accounts were distributed as dividends:

Year 1 $10,000
Year 2   45,000
Year 3   90,000

Determine the dividends in arrears for preferred stock at the end of second year.

a. $30,000
b. $10,000
c. $25,000
d. $0

please give the steps too.
Answer #1

Answer : C) $25,000

Annual Preferred Dividend =20,000*100*2% = 40,000

Divined Paid to Preferred stock For First and 2nd Years

Years 1 = 10,000 , arrears = 40,000-10,000 = 30,000

Years 2 = 45,000,  =(40,000 +5000 arrears ), arrears = 30,000-5000 = 25,000 (Answer)

Total Dividend Paid for Two years = 10,000+45,000 = 55,000

Total Divdend due for Twi years = 40,000*2 = 80,000

Arrears = 80,000-55,000 =25,000 (Answer)